OpenStudy (sydthekid913):

The weather forecast calls for high temperatures between 69° and 75°. Write this as an absolute value inequality, with t as temperature.

OpenStudy (sydthekid913):

please help and I'll give you a medal!

OpenStudy (sydthekid913):

hurry am being timed! Thank you soooooo much

OpenStudy (anonymous):

\(69\leq t\leq 75\) would be one way

OpenStudy (sydthekid913):

THANK YOU!!!

OpenStudy (anonymous):

ooh absolute value the average of \(69\) and \(75\) is \(72\) you can try \[|t-72|\leq 3\]

OpenStudy (anonymous):

go with the second one if it wants absolute value inequality

OpenStudy (sydthekid913):

kk